# Models That Prove Their Own Correctness

### v1

@@ A.2 Reinforcement Learning from Verifier Feedback
3      4:foreach $r=1,\dots R$ do
@@
Lemma A.6.
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. For any transcript $(x,y,q_{1},\dots,a_{R})$ we let $\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})$ denote the indicator random variable which equals 1 if and only if $V$ accepts the transcript. For any model $P_{\theta}$, denoting by $\mathrm{ver}_{V,\mu}(\theta)$ the verifiability of $P_{\theta}$ (Definition 3.3), it holds that
$\nabla_{\theta}\mathrm{ver}_{V,\mu}(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\\ (q_{r},a_{r})_{r=1}^{R}\end{subarray}}\left[\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\vec{d}_{s}(\theta)\right]$
@@
$\mathrm{ver}_{V,\mu}(\theta)$ $\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\end{subarray}}\left[V^{P_{\theta}}(x,y)\ \mathrm{accepts}\right]$
@@
Combining Equations 11 and 12, by linearity of expectation we have that
$\nabla_{\theta}\mathrm{ver}_{V,\mu}(\theta)$ $=\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\nabla_{\theta}\Pr_{\pi\sim\mathcal{T}^{\theta}(x)}[\pi=\pi^{\ast}]$
@@
Appendix B Annotations
@@
Appendix C A simple proof system for the GCD
@@
Claim C.1 (Bézout’s identity [Bezout, 1779]).
@@
Any coefficients $z_{0},z_{1}$ satisfying the assumption of C.1 are known as Bézout coefficients for $(x_{0},x_{1})$. C.1 immediately gives our simple proof system: For input $x=(x_{0},x_{1})$ and alleged GCD $y$, the honest prover sends (alleged) Bézout coefficients $(z_{0},z_{1})$. The Verifier accepts if and only if $y=z_{0}\cdot x_{0}+z_{1}\cdot x_{1}$ and $y$ divides both $x_{0}$ and $x_{1}$.
In this proof system the Verifier does not need to make any query; to fit within Definition 3.2, we can have the verifier issue a dummy query. Furthermore, by C.1 it is complete and has soundness error $s=0$. Lastly, we note that the Verifier only needs to perform two multiplications, an addition, and two modulus operations; in that sense, verification is more efficient than computing the GCD in the Euclidean algorithm.
@@
3:Update $q\coloneqq(r_{0}//r_{1})$, where ///// / denotes integer division.
@@
We formalize this in the terminology of Appendix B by defining a Transcript Annotator and Answer Extractor. Note that, since our proof system consists only of one “answer” $z$ send from the prover to the verifier, the entire transcript $\pi$ is simply $z=(z_{0},z_{1})$. [^10] Such a proof system is known as an NP-proof system and defines the complexity class NP [Karp, 1972].
@@
Appendix D Experiment details
@@
We use Karpathy’s nanoGPT. [^12] https://github.com/karpathy/nanoGPT We use a 6.7M parameter architecture of $8$ layers, $8$ attention heads, and $256$ embedding dimensions. We optimized hyperparameters via a random hyperparameter search, arriving at learning rate $0.0007$, AdamW $\beta_{1}=0.733$ and $\beta 2=0.95$, $10\%$ learning rate decay factor, no dropout, gradient clipping at $2.0$, no warmup iterations, and $10\%$ weight decay.
@@ Appendix E Additional figures for Section 5
Figure 5: Full version of Figure 2, with all $T\in\{0,3,4,5,6,7\}$ possibilities for the annotation cutoff in Annotated Transcript Learning ($T=0$ means no annotation). For a zoom-in of the right end of the axis, see Figure 2.

### v2

@@ A.2 Reinforcement Learning from Verifier Feedback
3      4:foreach Round of interaction $r=1,\dots R$ do
@@
Lemma A.6 (RLVF gradient estimation).
Fix an input distribution $\mu$ over $\Sigma^{\ast}$ and a verifier $V$ with round complexity $R$ and answer length $L_{a}$. For any transcript $(x,y,q_{1},\dots,a_{R})$ we let $\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})$ denote the indicator random variable which equals 1 if and only if $V$ accepts the transcript. For any model $P_{\theta}$, denote by $\mathrm{ver}(\theta)$ the verifiability of $P_{\theta}$ with respect to $V$ and $\mu$ (Definition 3.4). For any $\theta$, if $\mathrm{ver}(\theta)$ is differentiable in $\theta$, then
$\nabla\mathrm{ver}(\theta)=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\\ (q_{r},a_{r})_{r=1}^{R}\end{subarray}}\left[\mathrm{Acc}_{V}(x,y,q_{1},\dots,a_{R})\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in[L_{a}]\end{subarray}}\vec{d}_{s}(\theta)\right]$
@@
$\mathrm{ver}(\theta)$ $\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\end{subarray}}\left[\left<V,P_{\theta}\right>(x,y)\ \mathrm{accepts}\right]$
@@
Combining Equations 9 and 10, by linearity of expectation we have that
$\nabla_{\theta}\mathrm{ver}(\theta)$ $=\sum_{\pi^{\ast}\in\Pi^{\ast}(x)}\nabla_{\theta}\Pr_{\pi\sim\mathcal{T}^{\theta}(x)}[\pi=\pi^{\ast}]$
@@
Appendix B Preliminaries on Stochastic Gradient Ascent
For convenience of the reader, we provide a description of Stochastic Gradient Ascent and quote a theorem on its convergence. We adapt the presentation in Shalev-Shwartz and Ben-David [2014], noting that they present Stochastic Gradient Descent in its more general form for non-differentiable unbounded functions.
Stochastic Gradient Ascent (SGA) is a fundamental technique in concave optimization. Given a concave function $f\colon\mathbb{R}^{d}\to[0,1]$, SGA starts at $w_{0}=\vec{0}\in\mathbb{R}^{d}$ and tries to maximize $f(w)$ by taking a series of “steps.” Than directly differentiating $f$, SGA instead relies on an estimation $\nabla f(w)$: in each iteration, SGA takes a step in a direction that estimates $\nabla f(w)$.
Definition B.1 (Gradient estimator).
Fix a differentiable function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ for some $d$. A gradient estimator for $f$ is a randomized mapping $D_{f}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ whose expectation is the gradient of $f$. That is, for all $w\in\mathbb{R}^{d}$,
$\operatorname*{\mathbb{E}}_{v\sim D_{f}(w)}[v]=\nabla f(w).$
Note that this is an equality between $d$-dimensional vectors.
Hyperparameters: Learning rate $\lambda>0$ and number of iterations $N\in\mathbb{N}$.
Input: A function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ to maximize and a gradient estimator $D_{f}$ for $f$.
Output: A vector $\bar{w}\in\mathbb{R}^{d}$.
1
1:Initialize $w_{0}\coloneqq\vec{0}\in\mathbb{R}^{d}$.
22:for $i=0,\dots,N-1$ do
3:Sample $v_{i}\sim D_{f}(w_{i-1})$.
3      4:Update $w_{i}\coloneqq w_{i-1}+\lambda\cdot v_{i}$.
5:Output $\bar{w}\coloneqq\frac{1}{N}\sum_{i\in[N]}w_{i}$.
Algorithm 3 Stochastic Gradient Ascent
Theorem 14.8 in Shalev-Shwartz and Ben-David [2014] implies the following fact.
Fact B.2.
Fix a concave $f\colon\mathbb{R}^{d}\to[0,1]$, a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$, and upper-bounds $B_{\mathrm{Norm}},B_{\mathrm{Lip}}>0$. Let
$w^{\ast}\in\operatorname*{\mathrm{argmax}}_{w:||w||<B_{\mathrm{Norm}}}f(w),$
and let $\bar{w}$ denote the output of Algorithm 3 run for $N$ iterations with learning rate
$\lambda=\frac{B_{\mathrm{Norm}}}{B_{\mathrm{Lip}}\sqrt{N}}.$
If at every iteration it holds that $||d_{i}||<B_{\mathrm{Lip}}$, then
$\operatorname*{\mathbb{E}}_{\bar{w}}\left[f(\bar{w})\right]\geq f(w^{\ast})-\frac{B_{\mathrm{Norm}}\cdot B_{\mathrm{Lip}}}{\sqrt{N}}.$
Appendix C Annotations
@@
Appendix D A simple proof system for the GCD
@@
Claim D.1 (Bézout’s identity [Bezout, 1779]).
@@
Any coefficients $z_{0},z_{1}$ satisfying the assumption of D.1 are known as Bézout coefficients for $(x_{0},x_{1})$. D.1 immediately gives our simple proof system: For input $x=(x_{0},x_{1})$ and alleged GCD $y$, the honest prover sends (alleged) Bézout coefficients $(z_{0},z_{1})$. The Verifier accepts if and only if $y=z_{0}\cdot x_{0}+z_{1}\cdot x_{1}$ and $y$ divides both $x_{0}$ and $x_{1}$.
In this proof system the Verifier does not need to make any query; to fit within Definition 3.2, we can have the verifier issue a dummy query. Furthermore, by D.1 it is complete and has soundness error $s=0$. Lastly, we note that the Verifier only needs to perform two multiplications, an addition, and two modulus operations; in that sense, verification is more efficient than computing the GCD in the Euclidean algorithm as required by Remark 3.3.
@@
3:Update $q\coloneqq\lfloor r_{0}/r_{1}\rfloor$.
@@
We formalize this in the terminology of Appendix C by defining a Transcript Annotator and Answer Extractor. Note that, since our proof system consists only of one “answer” $z$ send from the prover to the verifier, the entire transcript $\pi$ is simply $z=(z_{0},z_{1})$. Since the verification is deterministic, this means that the proof system is of an $\mathsf{NP}$ type (however, note that the search problem of finding the “$\mathsf{NP}$-witness” $z=(z_{0},z_{1})$ is in fact in $\mathsf{P}$).
@@
Appendix E Experiment details
@@
We use Karpathy’s nanoGPT. [^11] https://github.com/karpathy/nanoGPT. We use a 6.7M parameter architecture of $8$ layers, $8$ attention heads, and $256$ embedding dimensions. We optimized hyperparameters via a random hyperparameter search, arriving at learning rate $0.0007$, AdamW $\beta_{1}=0.733$ and $\beta 2=0.95$, $10\%$ learning rate decay factor, no dropout, gradient clipping at $2.0$, no warmup iterations, and $10\%$ weight decay.